OpenStudy (uscrnamc):

how does the graph g(x) = [x] - 3 differ from the graph of f(x) = [x] a. shifted up by 3 units b. shifted down by 3 units c. shifted left by 3 units d. shifted right by 3 units

OpenStudy (amorfide):

if you have f(x)=x then it is just a straight line graph through the origin  where the origin is the point (0,0), this is the point in which it crosses the x and y axis now if I have g(x)=x+1  this is now a line through the point (0,1) now a coordinate is written as (x,y) where x is the x ordinate, and y is the y ordinate so from f(x) we know it goes through (0,0), which means x=0, y=0 and from g(x)=x+1 we know it goes through (0,1), which means x=0, y=1 so we can see that the y value has gone up by 1 there fore it has shifted up by 1 unit so a general form is, if we have f(x)=x and g(x)=x+b, then g(x) is shifted b units. if b is positive it is moved up, if b is negative it is moved down so in your question you have g(x)=[x] and f(x)=[x]-3 try to solve it using this information

OpenStudy (uscrnamc):

so since it is -3 i wanna say it shifted down

OpenStudy (amorfide):

wait no

OpenStudy (amorfide):

my last line is incorrect

OpenStudy (amorfide):

so in your question you have g(x)=[x]-3 and f(x)=[x] I wrote it the wrong way around sorry

OpenStudy (uscrnamc):

you got it right it shifted down 3 units

OpenStudy (amorfide):

yeah the answer is correct, I was correcting my self hahaha sorry! you are correct, g(x) is shifted 3 units down from f(x)! well done :)